1: The frequency of a sound wave was measured to be 432 Hz. If you know that the speed of sound in air was 343 m/s, then determine the wavelength of the sound wave.

2:The wavelength of a sound wave was measured to be 1.482 m. If you know that the speed of sound in air was 343 m/s, then determine the frequency of the sound wave.

3: If the wavelength and frequency of a sound wave was measured to be 1.01 m and 323 Hz, then determine the speed of sound.
Answer #1

1) We know that, where 'V' is the velocity of sound , 'f' is the frequency of the sound and '' is the wavelength of sound. So putting the given values of frequency(f) and velocity(V) of sound and solving for wavelength() ,we get

  

=> = 343/432 = 0.794 m

2) Given, = 1.482m, V = 343m/s

So,

=> f = 343/1.482 = 231.44 Hz

3) Given, = 1.01m , f = 323 Hz

So,

=> V = 323 1.01 = 326.23 m/s